OpenStudy (anonymous):

Help! Pretty Pwease!? (O///_///O) I already answered it, but I'm not sure if the correct answer is B or C.

OpenStudy (anonymous):

x^2+y^2=r^2 and r=4

OpenStudy (anonymous):

nope its C

OpenStudy (anonymous):

got it ??

OpenStudy (anonymous):

its C cuz 4^2=16

OpenStudy (anonymous):

keep rocking ;)